OpenStudy (anonymous):

Evaluate 5x2 - 4 when x = 3. help i am new

OpenStudy (anonymous):

http://openstudy.com/study#

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@Lindsay_552

OpenStudy (anonymous):

hhhheeellllppp

OpenStudy (anonymous):

can anyone help

OpenStudy (emogirlatemoooocow):

41

OpenStudy (anonymous):

r u sure

OpenStudy (emogirlatemoooocow):

im pretty sure.

OpenStudy (anonymous):

k thxs got u a medal

OpenStudy (emogirlatemoooocow):

otay

OpenStudy (emogirlatemoooocow):

5(3*2)−4 =(5)(9)−4 =45−4 =41